# Modeling temporal hypergraphs

### v1

@@ III.2 Commonly used temporal hyperedge statistics
We can define statistics to assess first-order effects and homophily with respect to exogenous variables, such as age or gender. The definitions of such statistics is very similar to subset repetition and assortativity of order one. The only change in the definition is that we replace the prior degree by a function giving the values of the respective variable. For example, if we have a binary variable $female(i)$, taking the value one if and only if $i$ is female, we can assess the first-order effect (testing whether females participate in events at a higher rate) via the statistic (which we also denote as $female$ by a slight abuse of notation):
$female(I)=\sum_{i\in I}female(i)\enspace.$ (9)
@@
$homophily(I)=-\sum_{\{i,i^{\prime}\}\in\binom{I}{2}}|female(i)-female(i^{\prime})|\enspace.$ (10)
In the Les Misérables data (augmented by the $gender$ variable as described above) we find that females are neither significantly overrepresented nor underrepresented in events (compared to their prevalence in the set of all 80 characters), but that there is homophily with respect to the “female” variable (see Model 9 in Table 1 and also compare Table 5). This means that actors typically co-appear with other actors of the same gender, more than what would be expected by chance alone.
III.3 Event probabilities
@@
The mean estimated probability over all actual chapters is $5\%$, exceeding even the probability of random guessing of the one-person chapters (that is $1/80$). Ten of the 288 chapters have a probability larger than or equal to $0.5$. Figure 2 shows some chapters of remarkably high probabilities, especially towards the end of the novel, where the model has a longer history to build upon. The two chapters with the highest probabilities ($p>0.99$) are Chapter 1.3.8, featuring the repeated appearance of eight actors (FN, DA, FV, and ZE together with their unfaithful lovers FT, LI, BL, and FA) and Chapter 5.1.21, featuring the re-appearance of six “friends of the ABC” together with Marius. Another remarkable chapter, regarding estimated probabilities, is the last one (Chapter 5.9.5, $p>0.87$, in Position 4 of 288 when ranking chapters by their predicted probabilities) in which Jean Valjean, Marius, and Cosette are happily reunited (sadly just moments before Jean Valjean dies), further “filling” the triad composed of the three most prominent actors.
@@
IV Model variants and other data
@@
$rec.d(I,E_{<t})=\sum_{(t_{m},I_{m})\in E_{<t}\colon I\subseteq I_{m}}\exp\left((t_{m}-t)\frac{\log 2}{T_{1/2}}\right)\enspace,$
@@
In the model definition above we assumed that event times are strictly ordered. However, in many real-world hypergraphs (e. g., emails within an organization or coautoring), several or many events may have the same time stamps. Conveniently, the theory of time-to-event models provides efficient ways to handle “tied event times” [68, 69, 70, 71] and statistical software typically can deal with multiple events occurring at the same point in time. More important is to understand the underlying assumptions. In the case of simultaneous events, RHEM assume that events with the same time stamp occur conditionally independent of each other, given the history of earlier events. The validity of this assumption depends on the given application area and also on the granularity of the time variable. To assess the robustness of findings, one may create several strictly ordered event sequences in which simultaneous events are randomly permuted and then check whether findings are affected in a meaningful way.
@@
V Conclusion
In this paper we reviewed relational hyperevent models (RHEM) [43, 38] and illustrated their use to define tailored null distributions for temporal hypergraphs and to test patterns such as preferential attachment, subset repetition, triadic closure, assortative mixing, and homophily. We argue that RHEM provide a versatile model family that can simultaneously reproduce in expectation observed distributions of any given vector of temporal hyperedge statistics. We sketched that RHEM have already been applied to diverse empirical settings and have been adapted to several generalizations of hypergraphs, among others to directed, typed, weighed, and signed hyperedges and to multilayer (or multimode) hypergraphs. RHEM have been applied to small networks, e. g., [12], as well as to networks with more than a million nodes and hyperedges [45]. The eventnet [^1] https://github.com/juergenlerner/eventnet software [38] provides a free and open-source reference implementation to compute a large selection of temporal hyperedge statistics.
@@ V Conclusion
The main motivation for the given analysis of the Les Misérables data was to illustrate the model. However, it nevertheless gives further empirical support to the insight (compare [52]) that when quantifying the overrepresentation or underrepresentation of given hypergraph configurations, such as filled triads, one should not only control for observed node degrees but also for the frequencies of higher-order subsets, such as dyads contained in the respective triad (compare the discussion of the JV–MA–CO triad given above in Section III.2).
@@
Acknowledgements.
We acknowledge financial support: J.L. from the Deutsche Forschungsgemeinschaft (DFG) Grant number 321869138; M.-G.H. from 4P-CAN project, ID 101104432, HORIZON-MISS-2022-CANCER-01, HORIZON-RIA; M.P. from the Slovenian Research and Innovation Agency (Javna agencija za znanstvenoraziskovalno in inovacijsko dejavnost Republike Slovenije) (Grant No. P1-0403).
Appendix A More detailed analysis of the Les Misérables data
@@
A variation of subset repetition is exact repetition, which captures a tendency to exactly repeat the participant list of prior events. We can test this effect with the statistic $repetition(I,E_{<t})$, defined in Eq. (6). Table 4 demonstrates that the Les Misérables data provides evidence for exact repetition, in addition to subset repetition of order one, two, and three. Findings on triadic closure have been already discussed in the main text. Results on further model specifications are given in the second row of Table 3.
We can test a gender-specific activity effect via the statistic $female(I)$, defined in Eq. (9), and we can test homophily with respect to gender with the statistic $homophily(I)$, defined in Eq. (10). In the Les Misérables data, augmented by a $gender$ variable as described in the main text, there are 28 female among the 80 actors. Thus, if actors were selected into chapters disregarding their gender (and not taking into account any other effect), we would expect that the average rate of females in chapters is $28/80$, that is 35%. In Table 5 we can see that female are not significantly overrepresented or underrepresented compared to this baseline ratio, but that there is homophily with respect to the “female” variable. This means that actors typically co-appear with other actors of the same gender, more than what would be expected by chance alone.

### v2

@@ 3.2 Commonly used temporal hyperedge statistics
We can define statistics to assess first-order effects and homophily with respect to exogenous variables, such as age or gender. The definition of such statistics is very similar to subset repetition and assortativity of order one. The only change in the definition is that we replace the prior degree by a function giving the values of the respective variable. For example, if we have a binary variable $female(i)$, taking the value one if and only if $i$ is female, we can assess the first-order effect (testing whether females participate in events at a higher rate) via the statistic (which we also denote as femalefemale by a slight abuse of notation):
$female(I)=\sum_{i\in I}female(i)\kern 5.0pt.$ (9)
@@
$homophily(I)=-\sum_{\{i,i^{\prime}\}\in\binom{I}{2}}|female(i)-female(i^{\prime})|\kern 5.0pt.$ (10)
In the Les Misérables data (augmented by the gendergender variable as described above) we find that females are neither significantly overrepresented nor underrepresented in events (compared to their prevalence in the set of all 80 characters), but that there is homophily with respect to the “female” variable (see Model 9 in Table 1 and also compare Table 5 in the appendix). This means that actors typically co-appear with other actors of the same gender, more than what would be expected by chance alone.
3.3 Event probabilities
@@
The mean estimated probability over all actual chapters is $5\%$, exceeding even the probability of random guessing of the one-person chapters (that is $1/80$). Ten of the 288 chapters have a probability larger than or equal to $0.5$. Figure 2 shows some chapters of remarkably high probabilities, especially towards the end of the novel, where the model has a longer history to build upon. The two chapters with the highest probabilities ($p>0.99$) are Chapter 1.3.8, featuring the repeated appearance of eight actors (FN, DA, FV, and ZE together with their unfaithful lovers FT, LI, BL, and FA) and Chapter 5.1.21, featuring the re-appearance of six “friends of the ABC” together with Marius. Another remarkable chapter, regarding estimated probabilities, is the last one (Chapter 5.9.5, $p>0.87$, in Position 4 of 288 when ranking chapters by their predicted probabilities) in which Jean Valjean, Marius, and Cosette are happily reunited, further “filling” the triad composed of the three most prominent actors.
@@
4 Model variants and other data
@@
$rec.deg(I,E_{<t})=\sum_{(t_{m},I_{m})\in E_{<t}\colon I\subseteq I_{m}}\exp\left((t_{m}-t)\frac{\log 2}{T_{1/2}}\right)\kern 5.0pt,$
@@
In the model definition above we assumed that event times are strictly ordered. However, in many real-world hypergraphs (e. g., emails within an organization or coauthoring), several or many events may have the same time stamps. Conveniently, the theory of time-to-event models provides efficient ways to handle “tied event times” [69, 70, 71, 72] and statistical software typically can deal with multiple events occurring at the same point in time. More important is to understand the underlying assumptions. In the case of simultaneous events, RHEM assume that events with the same time stamp occur conditionally independent of each other, given the history of earlier events. The validity of this assumption depends on the given application area and also on the granularity of the time variable. To assess the robustness of findings, one may create several strictly ordered event sequences in which simultaneous events are randomly permuted and then check whether findings are affected in a meaningful way.
@@
5 Conclusion
In this paper we reviewed relational hyperevent models (RHEM) [44, 39] and illustrated their use to define tailored null distributions for temporal hypergraphs and to test patterns such as preferential attachment, subset repetition, triadic closure, assortative mixing, and homophily. We argue that RHEM provide a versatile model family that can simultaneously reproduce in expectation observed distributions of any given vector of temporal hyperedge statistics. We sketched that RHEM have already been applied to diverse empirical settings and have been adapted to several generalizations of hypergraphs, among others to directed, typed, weighted, and signed hyperedges and to multilayer (or multimode) hypergraphs. RHEM have been applied to small networks, e. g., [12], as well as to networks with more than a million nodes and hyperedges [46]. The eventnet [^2] https://github.com/juergenlerner/eventnet software [39] provides a free and open-source reference implementation to compute a large selection of temporal hyperedge statistics.
@@ 5 Conclusion
The main motivation for the given analysis of the Les Misérables data was to illustrate the model. However, it nevertheless gives further empirical support to the insight (compare [53]) that when quantifying the overrepresentation or underrepresentation of given hypergraph configurations, such as filled triads, one should not only control for observed node degrees but also for the frequencies of higher-order subsets, such as dyads contained in the respective triad (compare the discussion of the JV–MA–CO triad given above in Section 3.2).
@@
Funding:
We acknowledge financial support: J.L. from the Deutsche Forschungsgemeinschaft (DFG) Grant number 555455503; M.-G.H. from 4P-CAN project, ID 101104432, HORIZON-MISS-2022-CANCER-01, HORIZON-RIA; M.P. from the Slovenian Research and Innovation Agency (Javna agencija za znanstvenoraziskovalno in inovacijsko dejavnost Republike Slovenije) (Grant No. P1-0403).
Data availability:
For illustration we analyze in this paper the network of actor co-appearance in chapters of Victor Hugo’s novel Les Misérables, compiled by Donald Knuth [52]. For reproducibility, we provide the preprocessed data, augmented by the “femalefemale” variable (as described in Section 2.2), in the file jean_events.csv, linked from the eventnet Webpage  [^3] https://github.com/juergenlerner/eventnet/tree/master/data/les_miserables.
Appendix A Appendix: more detailed analysis of the Les Misérables data
@@
We can test a gender-specific activity effect via the statistic $female(I)$, defined in Eq. (9), and we can test homophily with respect to gender with the statistic $homophily(I)$, defined in Eq. (10). In the Les Misérables data, augmented by a gendergender variable as described in the main text, there are 28 female among the 80 actors. Thus, if actors were selected into chapters disregarding their gender (and not taking into account any other effect), we would expect that the average rate of females in chapters is $28/80$, that is 35%. In Table 5 we can see that female are not significantly overrepresented or underrepresented compared to this baseline ratio, but that there is homophily with respect to the “female” variable. This means that actors typically co-appear with other actors of the same gender, more than what would be expected by chance alone.